Please solve the following quadratic equations by factors (Please do not use the quadratic formula)
x+7/3 + 2/x =4

3/x-1 - 2/x+1=1

x+2/x-4 = 2x+1/x-2

Answer by stanbon(75887)

Please solve the following quadratic equations by factors (Please do not use the quadratic formula)
-----
Since you did not put in parentheses I will have to
guess where the numerators and denominators start and end.
------
x+7/3 + 2/x =4
(x+7)/3 + 2/x = 4
Multiply thru by 3x to get:
x(x+7) + 2*2 = 4(3x)
x^2 + 7x + 4 = 12x
x^2 - 5x + 4 = 0
(x-4)(x-1) = 0
x = 4 or x = 1
====================
Cheers,
Stan H.
====================
3/x-1 - 2/x+1=1
x+2/x-4 = 2x+1/x-2